At 15 deg C, a bare wheel has a diameter of 30 cm, and the inside diameter of a steel rim is 29.930 cm. To what temperature must the rim be heated so as to slip over the wheel? For this type of steel, 𝝰 = 1.1 x 10^-5 /deg C.


1
Expert's answer
2021-04-30T11:23:54-0400

The rim of initial diameter 29.930 cm must be heated so its diameter is less than 30 cm:


"D_{30}=D_0(1+\alpha T),\\\space\\
T=\frac 1\alpha\bigg(\frac{D_{30}}{D_0}-1\bigg)=212°\text C."